Disprove the following statements using counterexample. 

1) (𝑥 + 4)2 = 𝑥2 + 16  

2) 𝑥2 > x  

3) √𝑥2 = 𝑥  


1
Expert's answer
2021-02-24T07:41:42-0500

1) Let "x=1"


"(x+4)^2=(1+4)^2=25"

"x^2+16=1^2+16=17"

"25\not=17"

Hence "(x+4)^2\not=x^2+16" for "x=1"


2) Let "x=\dfrac{1}{2}"


"x^2=(\dfrac{1}{2})^2=\dfrac{1}{4}<\dfrac{1}{2}=x"

Hence "x^2<x" for "x=\dfrac{1}{2}"


3) Let "x=-1"


"\sqrt{x^2}=\sqrt{(-1)^2}=\sqrt{1}=1\not=-1=x"

Hence "\sqrt{x^2}\not=x" for "x=-1."